OpenStudy (anonymous):

Complete the standard form of the equation –4y = 3 – 2x

OpenStudy (anonymous):

@Godlovesme

OpenStudy (anonymous):

-4y=3-2x

OpenStudy (anonymous):

@WhiteFangIII

OpenStudy (anonymous):

@Gabylovesyou

OpenStudy (godlovesme):

add 2x to both sided and u will get ?

OpenStudy (godlovesme):



OpenStudy (anonymous):

@perl can you help me?

OpenStudy (anonymous):

@gabgurl

OpenStudy (anonymous):

someone help me!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

@Godlovesme

OpenStudy (anonymous):

@Phebe

OpenStudy (phebe):

im here

OpenStudy (anonymous):

k can u help me

OpenStudy (phebe):

Slope = 1.000/2.000 = 0.500 x-intercept = 3/2 = 1.50000 y-intercept = -3/4 = -0.75000